# An upper bound of the silhouette validation metric for clustering

### v2

@@ Abstract
The silhouette coefficient quantifies, for each observation, the balance between within-cluster cohesion and between-cluster separation, taking values in [1,1][-1,1]. The average silhouette width (ASW) is a widely used internal measure of clustering quality, with higher values indicating more cohesive and well-separated clusters. However, the dataset-specific maximum of ASW is typically unknown, and the standard upper limit of 1 is rarely attainable. In this work, we derive for each data point a sharp upper bound on its silhouette width and aggregate these to obtain a canonical upper bound on the ASW. This bound—often substantially below 1—enhances the interpretability of empirical ASW values by indicating how close a given clustering result is to the best possible outcome on that dataset. It can be used to confirm global optimality, guide the evaluation of clustering solutions, and be refined to incorporate minimum cluster-size constraints for greater practical relevance. Finally, we extend the framework to establish a corresponding bound for the macro-averaged silhouette.
@@ 1 Introduction
Cluster analysis is a fundamental tool in data science for discovering structure in unlabeled data [aggarwal2013introduction]. Its applications span the natural and social sciences, engineering, education, economics, and the health sciences [dinh2025data]. Because ground-truth labels are rarely available, evaluating clustering quality is challenging, so researchers rely on internal validation indices that summarize within-cluster compactness and between-cluster separation [jain1999data].
@@
Among the cluster validation indices, the silhouette width (also known as the silhouette coefficient, silhouette score, or silhouette index) is one of the most widely used measures to assess how well each point fits within its assigned cluster relative to its nearest neighboring cluster [ogsilhouette]. The average of this measure across all data points is often referred to as the average silhouette width $\mathrm{ASW}$. $\mathrm{ASW}$ values close to 1 signal well‑separated, compact clusters; values near 0 imply overlapping clusters; and negative values indicate that points are likely misassigned. Extensive work has examined and refined $\mathrm{ASW}$. [batool2021clustering] proved that it satisfies key axioms for clustering quality measures such as scale invariance and consistency. More recently, [john] examined alternative aggregation strategies that improve $\mathrm{ASW}$’s robustness to cluster-size imbalance.
Although $\mathrm{ASW}$ is widely used, its raw values are difficult to interpret across heterogeneous datasets. Intrinsic data characteristics impose strict ceilings on attainable $\mathrm{ASW}$ scores; for example, when clusters overlap or deviate from convex shapes, even the best partition may yield a low $\mathrm{ASW}$. Comparing $\mathrm{ASW}$ values across datasets with different properties can therefore lead to misguided conclusions about cluster quality.
To address this issue, we derive a data-dependent upper bound on the $\mathrm{ASW}$. This bound supplies essential context: if analysis shows that no partition can exceed, say, an $\mathrm{ASW}$ of 0.30, an empirical result of 0.29 should be regarded as essentially optimal. Such insight prevents wasted effort in pursuit of negligible gains and enables more informed evaluation of clustering quality.
@@
The biggest obstacle in silhouette optimization is the combinatorial explosion of the solution space as the number of samples increases. Even for moderately sized datasets, the number of possible clusterings is astronomical, rendering optimization through exhaustive search with respect to any internal criterion computationally prohibitive. [batool2021clustering] addressed this limitation by proposing two heuristic algorithms that search for local maxima of $\mathrm{ASW}$, noting that global optimization is tractable only for the smallest datasets. Our contribution complements their approach by providing a data-aware benchmark against which the quality of such locally optimal solutions can be evaluated.
@@
The main contributions are as follows:
    We introduce a novel data–dependent upper bound on the $\mathrm{ASW}$, computable in $O(n^{2}\!\log n)$ time, to yield a global ceiling that no clustering of the given dissimilarity matrix can surpass.
    We evaluate the proposed upper bound on large-scale synthetic datasets and a diverse suite of real-world datasets across multiple dissimilarity measures. We show that the bound is nearly tight in many of the cases examined.
The rest of this paper is organized as follows. Section 2 reviews related works. Section 3 introduces preliminaries. Section 4 details the proposed method. Section 5 presents the experimental results. Finally, section 6 summarizes the findings and outlines directions for future research.
2 Related work
Silhouette coefficientEnergy & Power Systems[ouyang2024k][rodriguez2024novel][zheng2025practical][duarte2021increasing][eskandarnia2022embedded][zhao2023generalized][koutsandreas2024harnessing][zhou2025recognition][hu2023transformer][alrasheed2024building][diaz2022typologies]Transportation & Mobility[mohd2024passengers][wang2021analyzing][bai2023adaptive][sheng2025multi][guo2025adaptive][passarella2024anomaly][ma2025unravelling][li2025clustering][ansari2021spatiotemporal]Biomedical & Health[jamalpour2025high][sharafoddini2021identifying][ghaderi2023self][gill2024symptom][baier2024nocturnal][kim2022pathological][rustam2022data][jha2021apache][xian2025systematic][bhagat2022dpcf]Food & Agriculture[edris2024identifying][singh2025development][gere2023recommendations]Social & Economic[owoo2025hierarchical][quintero2022analysis][gulma2022new][zhang2025two][aonso2024cluster]Materials & Chemistry[pham2025clustering][bi2025comprehensive][sancho2022cluster][bagirov2023novel][consuegra101unsupervised]Environment & Earth Sciences[nourani2022robust][guo2023combining][yang2022zoning][huang2023dominant][balbi2024hierarchical][biabiany2020design][guan2025novel][gupta2025criteria][zhao2023comparison][deng2025on]Sensor Networks & IoT[kotary2020distributed][kotary2021many][panwar2024distributed][panwar2025distributed][waidyanatha2023regularity]Big Data[qaroush2025maximum][thirumoorthy2021hybrid][song2023novel][chander2023dolphin]Cluster Validity & Selection[khan2025standardizing][khan2025standardization][batool2021clustering][lenssen2024medoid][capo2023fast][gagolewski2021are][bagirov2023finding]Algorithm Development[koren2025preprocessing][silva2025automatic][manochandar2020development][qin2025single][zhang2025two][bai2023adaptive][sheng2025multi][kotary2020distributed][kotary2021many][du2024multi][fu2025gsscl][gan2024discrimination][fu2023gc][enriquez2025performance][gere2023recommendations]Feature & Dim‑Reduction[rustam2022data][dwivedi2025novel][zhou2024ocae][chander2023dolphin]
Figure 2: Taxonomy of how the silhouette coefficient is employed across research fields.
@@ 2 Related work
The silhouette coefficient [ogsilhouette] is a distance-based internal validity index that quantifies how well each object fits into its assigned cluster relative to its nearest alternative. Fig. 1 illustrates the construction. For a point $x_{i}$ in cluster $C_{\ell}$, we measure within-cluster cohesion by $a(x_{i})$—its average dissimilarity to the other members of $C_{\ell}$—and nearest-cluster separation by $b(x_{i})$—its average dissimilarity to the closest neighboring cluster. These quantities yield an individual silhouette score $s(x_{i})\in[-1,1]$, which is large when separation dominates cohesion, near zero around decision boundaries, and negative when $x_{i}$ is, on average, closer to another cluster than to its own. Aggregating over points gives the ASW, a dataset-level summary of partition quality.
In practice, ASW is often used for model selection by choosing the number of clusters $K$ that maximizes it. Despite numerous extensions—e.g., to categorical data [dinh2019estimating], fuzzy clustering [lasek2024employing], hierarchical clustering [owoo2025hierarchical], model-based clustering [guan2025novel], spectral clustering [xian2025systematic], spherical clustering [deng2025on], and optimization-based methods [yang2024clustering]—the original silhouette framework remains a standard baseline for benchmarking clustering algorithms.
@@
Other commonly used internal validation indices include the Davies–Bouldin index [dbindex] and the Calinski–Harabasz index [chindex], which also evaluate compactness and separation but rely on alternative formulations of intra- and inter-cluster dispersion.
@@
In this work, we focus on the silhouette due to its interpretability and widespread adoption across both research and applied domains. To account for cluster-size imbalance, we additionally employ the macro-averaged silhouette, defined as the unweighted mean of per-cluster silhouette averages. This variant ensures that smaller clusters contribute equally to the overall assessment, providing a more balanced view of clustering quality.
@@
PAMSIL (Partitioning Around Medoids based on SILhouette) is a modification of the classical PAM algorithm [kaufman90], designed to directly optimize the ASW [pamsil]. Unlike $K$-means or standard $K$-medoids, which minimize within-cluster dissimilarities, PAMSIL explicitly seeks cluster configurations that maximize ASW. It achieves this by modifying the SWAP phase of PAM: in each iteration, the algorithm performs the swap that yields the largest improvement in ASW and terminates once no further improvement is possible, ensuring convergence to a local optimum. An efficient implementation of PAMSIL is available in the open-source K-Medoids Python package [Schubert2022]. Moreover, Fatima Batool and Christian Hennig have proposed alternative algorithms that find local ASW optima [batool2021clustering].
@@
When ground-truth class labels are available, external validation measures provide a way to quantify how well a clustering agrees with a known reference partition. Two of the most commonly used indices are the adjusted Rand index (ARI) [ari] and the adjusted mutual information (AMI) [ami]. Both adjust their underlying similarity measures for chance agreement, ensuring that random labelings yield expected scores near zero.
The ARI extends the classical Rand index by comparing all pairs of samples and counting how consistently they are assigned to the same or different clusters in both partitions. Its normalization accounts for random overlap, producing values in [1,1][-1,1] with $1$ indicating identical partitions. The AMI applies a similar correction to the mutual information (MI) between partitions, normalizing it by the expected MI under random assignments and bounding the result between 0 and $1$.
Both indices are symmetric, easy to interpret, and widely used in benchmarking studies, where they enable consistent comparison of clustering algorithms across datasets with known class structure.
@@ 3 Preliminaries
Figure 3: The first 27 Bell numbers, which enumerate the partitions of an $n$-element set. This rapid growth highlights the combinatorial explosion of clustering possibilities as $n$ increases.
For our purposes, it will be convenient to work with the off-diagonal entries of $\Delta$, sorted row-wise in ascending order. We denote this transformed dissimilarity matrix by $\hat{\Delta}$ ($\hat{\Delta}\in\mathbb{R}^{n\times n-1}$). In other words, we have for every $i$ $\{\Delta_{ij}\}_{j=1}^{n}\setminus\{\Delta_{ii}\}=\{\hat{\Delta}_{ij}\}_{j=1}^{n-1}$ and $\hat{\Delta}_{i1}\leq\hat{\Delta}_{i2}\leq\cdots\leq\hat{\Delta}_{i(n-1)}$. Constructing $\hat{\Delta}$ requires sorting each of the $n$ rows of $\Delta$, which costs $\mathcal{O}(n^{2}\log n)$ time in total.
@@
Let $K\in\{2,\ldots,n\}$. A $\mathbf{K}$-clustering of $\{i\}_{i=1}^{n}$ is a set $\mathcal{C}_{K}=\{C_{I}\}_{I=1}^{K}$ such that $C_{I}\subsetneq\{i\}_{i=1}^{n}$ for any $I$, and
@@
The total number of clusterings of an $n$-element set grows according to the Bell numbers (Fig. 3), illustrating the combinatorial complexity of the clustering problem.
We will now define the silhouette width, which was introduced in 1987 by P.J. Rousseeuw [ogsilhouette]. We employ similar notation.
@@
Let $\mathcal{C}_{K}$ be a $K$-clustering of the dataset $\{1,\dots,n\}$ with dissimilarity matrix $\Delta$. The silhouette width for data point $i\in C_{I}$ is 0 if $|C_{I}|=1$; otherwise it is
$s(i|\mathcal{C}_{K},\Delta)\coloneqq\frac{b(i)-a(i)}{\max\{a(i),b(i)\}},$
@@
The average silhouette width (ASW) is defined as
$ASW(\mathcal{C}_{K},\Delta)\coloneqq\frac{1}{n}\sum_{i=1}^{n}s(i|\mathcal{C}_{K},\Delta).$
@@
Individual silhouette widths provide a valuable graphical diagnostic, but for large $n$ it is customary to summarize the clustering quality by $\operatorname{ASW}$ [medoid]. A common heuristic is to run a clustering algorithm for several choices of $K$ and select the value that maximizes $\operatorname{ASW}$. When even that maximum is low, it is possible that the limitation stems from the intrinsic structure of the dissimilarity data. This raises the natural question: given a dissimilarity matrix, what is the highest possible $\mathrm{ASW}$? To discuss this question, let $\mathcal{C}^{*}$ denote a silhouette-optimal clustering.
Definition 4 (Silhouette-optimal clustering).
Let $\Delta$ be the dissimilarity matrix of the dataset $\{1,\ldots,n\}$. We define
$\mathfrak{C}\coloneqq\{\mathcal{C}_{K}:\mathcal{C}_{K}\text{ is a $K$-clustering of $\{1,\ldots,n\}$}\}$
to be the set of all possible clusterings. Any
$\mathcal{C}^{*}\in\operatorname*{argmax}_{\mathcal{C}_{K}\in\mathfrak{C}}\{\mathrm{ASW}(\mathcal{C}_{K},\Delta)\}$
is called a silhouette-optimal clustering.

### v4

@@ Abstract
The silhouette coefficient quantifies, for each observation, the balance between within-cluster cohesion and between-cluster separation, taking values in the range [1,1][-1,1]. The average silhouette width ($\operatorname{ASW}$) is a widely used internal measure of clustering quality, with higher values indicating more cohesive and well-separated clusters. However, the dataset-specific maximum of $\operatorname{ASW}$ is typically unknown, and the standard upper limit of 1 is rarely attainable. In this work, we derive for each data point a sharp upper bound on its silhouette width and aggregate these to obtain a canonical upper bound of the $\operatorname{ASW}$. This bound—often substantially below 1—enhances the interpretability of empirical $\operatorname{ASW}$ values by providing guidance on how close a given clustering result is to the best possible outcome for that dataset. We evaluate the usefulness of the upper bound on a variety of datasets and conclude that it can meaningfully enrich cluster quality evaluation; however, its practical relevance depends on the specific dataset. Finally, we extend the framework to establish an upper bound of the macro-averaged silhouette.
@@ 1 Introduction
Cluster analysis is a fundamental tool in data science for discovering structure in unlabeled data [1]. Its applications span the natural and social sciences, engineering, education, economics, and the health sciences [9]. Because ground truth labels are rarely available, evaluating clustering quality is challenging, so researchers rely on internal validation indices that summarize within-cluster compactness and between-cluster separation [11].
@@
Among the internal cluster validation indices, the silhouette width (also known as the silhouette coefficient, silhouette score, or silhouette index) is one of the most widely used measures. It quantifies how well each point fits within its assigned cluster relative to its nearest neighboring cluster [18]. The average of this measure across all data points is often referred to as the average silhouette width ($\operatorname{ASW}$). Silhouette scores close to 1 signal well-separated, compact clusters; values near 0 typically occur for points that lie between two clusters; a negative value indicates that a point is on average closer to its nearest neighboring cluster than to its own and is therefore commonly used to flag a potential misplacement. These interpretations are most reliable when clusters are roughly isotropic and of comparable scale; when cluster diameters differ substantially, silhouette values can be misleading, a limitation we return to in Section 6.
Extensive work has examined and refined $\operatorname{ASW}$. Batool et al. [4] proved that it satisfies key axioms for clustering quality measures such as scale invariance and consistency. More recently, Pavlopoulos et al. [16] examined alternative aggregation strategies that improve $\operatorname{ASW}$’s robustness to cluster size imbalance. The so-called medoid-based silhouette is a simple approximation of the silhouette coefficient that replaces average distances to clusters with distances to cluster medoids, making computations cheaper [14]. The medoid of a cluster is the cluster member with the smallest total distance to all other members [20].
Although $\operatorname{ASW}$ is widely used, its raw values are difficult to interpret. Intrinsic data characteristics impose strict ceilings on attainable $\operatorname{ASW}$ scores; for example, when clusters overlap or deviate from convex shapes, even the best partition may yield a low $\operatorname{ASW}$. As a result, observing a low $\operatorname{ASW}$ value alone does not reveal whether this reflects poor clustering performance or an inherent limitation of the data itself. For any dataset with more than a few samples, the true maximum ASW is unknown, which makes it difficult to assess clustering performance in relative terms. Relating an empirical $\mathrm{ASW}$ to a data-dependent upper bound could mitigate this problem by indicating how close the given solution is to the best possible outcome.
Against this background, we pose the following research question.
Question: Given a dissimilarity matrix, can we efficiently compute an upper bound of the $\operatorname{ASW}$ that provides meaningful guidance on how close an empirical clustering is to the global $\operatorname{ASW}$ maximum?
In connection with this question, our main contributions are as follows:
    We introduce a novel data-dependent upper bound on the $\operatorname{ASW}$, computable in $O(n^{2}\!\log n)$ time, to yield a global ceiling that no clustering of the given dissimilarity matrix can surpass.
    We evaluate the usefulness of the proposed upper bound on synthetic datasets and a diverse suite of real-world datasets across multiple dissimilarity measures.
In theory, the proposed upper bound can provide important context for interpretation: if analysis shows that no partition can exceed, say, an $\operatorname{ASW}$ of $0.30$, then an empirical result of $0.29$ can be understood as being close to the best achievable value for the given data. Such information helps practitioners assess how much room for improvement remains and avoid expending effort in search of gains that are provably limited by the structure of the data.
@@
The biggest obstacle in silhouette optimization is the combinatorial explosion of the solution space as the number of samples increases. Even for moderately sized datasets, the number of possible clusterings is astronomical, rendering optimization through exhaustive search with respect to any internal criterion computationally prohibitive. Batool et al. [4] addressed this limitation by proposing two heuristic algorithms that search for local maxima of $\operatorname{ASW}$, noting that global optimization is tractable only for the smallest datasets. Our contribution complements their approach by providing a data-aware benchmark against which the quality of such locally optimal solutions can be evaluated. Figure 1 illustrates the intended workflow.
1. Compute upper bound2. Measure empirical ASW3. Estimate proximity toglobal optimality
Figure 1: Illustrative workflow for using the upper bound as a diagnostic tool. The process involves (1) computing the $\mathrm{ASW}$ upper bound from the given dissimilarity matrix, (2) measuring the $\mathrm{ASW}$ achieved by the chosen clustering algorithm, and (3) assessing the proximity to global optimality by analyzing the gap between these two values.
To summarize: this paper’s main objective is to propose and assess a method for computing an upper bound of the $\mathrm{ASW}$ for a given dissimilarity matrix. We believe this contribution is important because such a data-dependent upper bound could be useful in clustering quality evaluations.
The rest of this paper is organized as follows. Section 2 reviews related works. Section 3 introduces preliminaries. Section 4 details the proposed method. Section 5 presents and discusses the experimental results. Finally, Section 6 summarizes the findings and outlines directions for future research.
Lastly, Table 1 summarizes the notation used throughout the paper.
Table 1: Summary of notation used throughout the paper.
Symbol Description
$n$ Number of objects in the dataset
$i,j$ Indices of objects, in $\{1,\dots,n\}$
$\Delta\in\mathbb{R}^{n\times n}$ Dissimilarity matrix with entries $\Delta_{ij}$
$\hat{\Delta}\in\mathbb{R}^{n\times(n-1)}$ Row-wise sorted off-diagonal dissimilarities
$K$ Number of clusters
$\mathcal{C}_{K}=\{C_{I}\}_{I=1}^{K}$ A $K$-clustering of $\{1,\dots,n\}$
$C_{I}$ Cluster $I$; $|C_{I}|$ denotes its size
$\mathcal{C}^{*}$ Silhouette-optimal clustering
$a(i)$ Average dissimilarity from object $i$ to members of its own cluster
$b(i)$ Minimum average dissimilarity from $i$ to points in any other cluster
$s(i\mid\mathcal{C}_{K},\Delta)$ Silhouette width of object $i$
$\operatorname{ASW}(\mathcal{C}_{K},\Delta)$ Average silhouette width
$S_{\mathrm{macro}}(\mathcal{C}_{K},\Delta)$ Macro-averaged silhouette
$q(i,\Delta,k)$ $k$-quotient of object $i$ used in the upper bound construction
$m$ Minimum allowed cluster size in constrained solution spaces
$f(i,\Delta)$, $f_{m}(i,\Delta)$ Minima of $q(i,\Delta,k)$ over $k$ (unconstrained and with $\min_{I}|C_{I}|\geq m$)
$\texttt{UB}(\Delta)$ Global upper bound of $\operatorname{ASW}$
$\texttt{UB}_{m}(\Delta)$ Upper bound of $\operatorname{ASW}$ with minimum cluster size $m$
2 Related work
@@
The silhouette coefficient [18] is a distance-based internal validity index that quantifies how well each object fits into its assigned cluster relative to its nearest alternative. Figure 2 illustrates the construction. For a point $x_{i}$ in cluster $C_{I}$, we measure within-cluster cohesion by $a(x_{i})$—its average dissimilarity to the other members of $C_{I}$—and nearest-cluster separation by $b(x_{i})$—its average dissimilarity to the closest neighboring cluster. These quantities yield an individual silhouette score $s(x_{i})\in[-1,1]$, which is large when separation dominates cohesion, near zero around decision boundaries, and negative when $x_{i}$ is, on average, closer to another cluster than to its own. Aggregating over points gives the $\operatorname{ASW}$, a dataset-level summary of partition quality.
@@
In practice, $\operatorname{ASW}$ is often used for model selection by choosing the number of clusters $K$ that maximizes it [2, 25, 3]. Other commonly used internal validation indices include the Davies–Bouldin index [7] and the Calinski–Harabasz index [5], which also evaluate compactness and separation but rely on alternative formulations of intra- and inter-cluster dispersion.
@@
In this work, we focus on the silhouette due to its interpretability and widespread adoption across both research and applied domains. To account for cluster size imbalance, we additionally employ the macro-averaged silhouette [16], defined as the unweighted mean of per-cluster silhouette averages. This variant ensures that smaller clusters contribute equally to the overall assessment, providing a more balanced view of clustering quality.
@@
PAMSIL (Partitioning Around Medoids based on SILhouette) is a modification of the classical PAM algorithm [12], designed to directly optimize the $\operatorname{ASW}$ [22]. Unlike $K$-means or standard $K$-medoids, which minimize within-cluster dissimilarities, PAMSIL explicitly seeks cluster configurations that maximize $\operatorname{ASW}$. It achieves this by modifying the SWAP phase of PAM: in each iteration, the algorithm performs the swap that yields the largest improvement in $\operatorname{ASW}$ and terminates once no further improvement is possible, ensuring convergence to a local optimum. An efficient implementation of PAMSIL is available in the open-source kmedoids Python package [19]. Moreover, Fatima Batool and Christian Hennig have proposed alternative algorithms that find local $\operatorname{ASW}$ optima [4].
In practice, PAMSIL can be computationally demanding, particularly for larger datasets. For this reason, accelerated variants have been developed. The kmedoids package includes FastMSC and DynMSC, which are both methods for optimizing the medoid-based version of the silhouette. DynMSC automatically explores different values of $K$ and returns the clustering with the highest average medoid silhouette.
Despite these developments around the silhouette coefficient, the existing literature largely focuses on 1) comparing the silhouette score with other clustering quality indices [6], 2) proposing closely related alternatives that are preferable in certain contexts [16, 22], or 3) designing algorithms for silhouette-optimization [22, 4, 14, 19]. To the best of our knowledge, these works do not provide an efficiently computable, data-dependent upper bound on the best achievable $\mathrm{ASW}$ for a given dataset. Our contribution complements these lines of work by deriving such a bound directly from the pairwise dissimilarities, enabling empirical silhouette scores to be interpreted relative to a dataset-specific ceiling and providing a principled benchmark for silhouette-optimization heuristics.
@@
When ground truth class labels are available, external validation measures provide a way to quantify how well a clustering agrees with a known reference partition. Two of the most commonly used indices are the adjusted Rand index (ARI) [10] and the adjusted mutual information (AMI) [23]. Both adjust their underlying similarity measures for chance agreement, ensuring that random labelings yield expected scores near zero.
The ARI extends the classical Rand index, which compares all pairs of samples and counts how consistently they are assigned to the same or different clusters in both partitions. Its normalization accounts for random overlap and produces values in [1,1][-1,1] with $1$ indicating identical partitions. The AMI applies a similar correction to the mutual information (MI) between partitions, normalizing it by the expected MI under random assignments and bounding the result between 0 and $1$.
Both indices are symmetric and widely used in benchmarking studies, where they enable consistent comparison of clustering algorithms across datasets with known class structure.
@@ 3 Preliminaries
Let $K$ be an integer between $1$ and $n$. A $\mathbf{K}$-clustering of $\{i\}_{i=1}^{n}$ is a set $\mathcal{C}_{K}=\{C_{I}\}_{I=1}^{K}$ such that $C_{I}\subseteq\{1,\ldots,n\}$ for all $I$, and
@@
Remark. The case $K=1$ corresponds to the trivial clustering $\mathcal{C}_{1}=\{\{1,\ldots,n\}\}$, while $K=n$ corresponds to the trivial partition into singletons.
In the continuation of the article, we only consider $K>1$ for convenience. Although we emphasize that the method proposed in Section 4 does not require such an assumption, and can be applied to datasets with a single cluster as well.
The total number of clusterings of an $n$-element set grows according to the Bell numbers $B_{n}$. For instance, the number of possible partitions of only 26 objects is $B_{26}=49631246523618756274$.
We will now define the silhouette width, which was introduced in 1987 by P.J. Rousseeuw [18]. We employ similar notation.
@@
Let $\mathcal{C}_{K}$ be a $K$-clustering of the dataset $\{1,\dots,n\}$ with dissimilarity matrix $\Delta$. The silhouette width for data point $i\in C_{I}$ is defined as
$s(i\mid\mathcal{C}_{K},\Delta)\coloneqq\begin{cases}\dfrac{b(i)-a(i)}{\max\{a(i),b(i)\}}&\text{ if }|C_{I}|>1\land\max\{a(i),b(i)\}>0,\\[3.00003pt] 0&\text{ if }\left(|C_{I}|>1\land\max\{a(i),b(i)\}=0\right)\lor|C_{I}|=1\end{cases}$
@@
The average silhouette width ($\operatorname{ASW}$) is defined as
$\operatorname{ASW}(\mathcal{C}_{K},\Delta)\coloneqq\frac{1}{n}\sum_{i=1}^{n}s(i\mid\mathcal{C}_{K},\Delta).$
@@
Individual silhouette widths provide a useful graphical diagnostic, but for large $n$ it is customary to summarize the clustering quality by $\operatorname{ASW}$ [14]. A common heuristic is to run a clustering algorithm for several choices of $K$ and select the value that maximizes $\operatorname{ASW}$. When even that maximum is low, it is possible that the limitation stems from the intrinsic structure of the dissimilarity data. This raises the natural question: given a dissimilarity matrix, what is the highest possible $\operatorname{ASW}$?
In the remainder of the paper, we let $\mathcal{C}^{*}$ denote a silhouette-optimal clustering. That is, if $\mathcal{C}_{K}$ is a $K$-clustering of $\Delta$ then $\mathrm{ASW}(\mathcal{C}_{K},\Delta)\leq\mathrm{ASW}(\mathcal{C}^{*},\Delta)$.